OpenStudy (anonymous):

solve. 15x^4-28x^2+5=0

OpenStudy (anonymous):

x=1/sqrt(5) x=-1/sqrt(5) and x=sqrt(5/3) and x=-sqrt(5/3)

OpenStudy (anonymous):

(5x^2-1)(3x^2-5) \[x= \pm \sqrt{1/5}and \pm \sqrt{5/3}\]